THE OPTICAL STARK SPECTRA OF CoF AND CoH

HAILING WANG, \underline{XIUJUAN ZHUANG}, AND TIMOTHY C. STEIMLE , Department of Chemistry and Biochemistry, Arizona State University, Tempe,AZ 85287.

Comparing the predicted and observed permanent electric dipole moments, μ_{el}, for simple diatomic cobalt containing molecules is a means of testing computational methodologies used for modeling cobalt-catalyzed reactions. Here we report on the experimental determination of μ_{el} for CoH and CoF from the analysis of the Stark shifts in the (0,0) bands of the A′^{3}Φ_{4} − X^{3}Φ_{4} system of cobalt monohydride, CoH, and the [18.8]^{3}Φ_{4} − X^{3}Φ_{4} system of cobalt monofluoride, CoF. In addition, the proton magnetic hyperfine interaction in CoH is analyzed. The relative values for μ_{el} are rationalized using a molecular orbital correlation diagram and with high-level ab initio predictions^{a}.

aTomonari,M.; Okuda, R.; Nagashima, U.; Tanaka, K.; Hirano, T.;J. Chem. Phys. \underline{126} 144307, 2007.